OpenStudy (anonymous):

Write the equations of these lines in slope-intercept form 1. The slope is 3/2; the y-intercept is 3. 2. The slope is -1/4; the y-intercept is 1. 3. The slope is -6; the y-intercept is 3/4 4. The slope is -3; the y-intercept is -1/2 5.6.

OpenStudy (lxelle):

1. y= 3/2x+3 Using the y=mx+c formula. The slope which is gradient is m. Y intercept is c.